Ifa^x=b^y=c^z , xyz=1 then x+y+z=?Please give the solution

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Dev
There is some modifications in question.
a^(1/x)=b^(1/y)=c^(1/z)=k(say). then a=k^x, b=k^y, c=k^z. Given abc=1 therefore k^(x +y +z)=1=k^0 this imply x +y+ z=0
Regards
Arun (askIITians forum expert)